OpenStudy (anonymous):

Find S8 for the geometric series 3 + -6 + 12 + -24 +… A. -384 B. 192 C. -96 D. -255

OpenStudy (johnweldon1993):

\[S_n = a \frac{ 1 - r^n }{ 1 - r }\] 'a' represents the first term you have (3) 'n' represents the amount of numbers you want to sum (8) 'r' represents the common ratio....(?) what is your common ratio?

OpenStudy (anonymous):

n=8 ? a = -3 what does r = ?

OpenStudy (johnweldon1993):

right n = 8 a = 3 not -3 right? r...what do we multiply 3 by..to get -6? and what do we multiply -6 by....to get 12?

OpenStudy (anonymous):

-2

OpenStudy (johnweldon1993):

perfect so put everything together knowing a = 3 r = -2 and n = 8 \[S_8 = 3\frac{ 1 + 2^8 }{ 1 + 2 }\]

OpenStudy (johnweldon1993):

Wait a minute...is this finding a sum? or finding the 8th term??

OpenStudy (anonymous):

i beilve the sum

OpenStudy (johnweldon1993):

It must be finding the 8th term....because the equation doesn't work for me and when finding the 8th term...I get -384 which is one of the answers. let me recheck real quick

OpenStudy (johnweldon1993):

nvm wrote it wrong..ignore the above equation post

OpenStudy (johnweldon1993):

\[S_8 = 3\frac{ 1 - (-2^8) }{ 1 + 2 }\] forgot those parenthesis and the fact that -2 gets the exponent first

OpenStudy (johnweldon1993):

and this does indeed give me one of those answers...sorry about that confusion

OpenStudy (johnweldon1993):

And once you do that CORRECT equation *again sorry about the confusion* you should get -255